# Xing Huang(a), Eva Kopfer(b), Pierre Monmarché(c), Panpan Ren(d) a)

### v1

@@ 1 Introduction
The first exponential convergence result stated in terms of relative entropy has been obtained in the seminal work [58] of Villani. It introduced the notion of hypocoercivity which has been developped since then in various directions, see [18, 44, 1, 10, 6, 8] and references therein on this general topic. More specifically, [58, Theorem 39] gives an exponential decay in relative entropy along (1.3) provided $V\in C^{2}(\mathbb{R}^{d})$ and has a bounded Hessian, and $\mu$ satisfies a so-called log-Sobolev inequality (see (3.3) below) with respect to the Dirichlet form $\mathscr{E}(f,g)=\mu(\nabla f\cdot\nabla g)$, where $\nabla$ is the gradient on $\mathbb{R}^{2d}$.
@@
However, the result in [58, Theorem 39], along with nearly all subsequent works, relies on having an explicit representation of the invariant probability measure $\mu$. An exception arises in the so-called uniformly dissipative case, which is a very restrictive situation where the drift of the SDE induces a deterministic contraction along the stochastic flow [5, 40, 42]. In particular, there are many tools to establish log-Sobolev inequalities [3], but most of them require an explicit expression, or are restricted to the uniformly-dissipative case as the Bakry-Émery criterion (which in its classical form is written for explicit invariant measures but in fact apply more generally [42]).
@@ 2.1 Exponential Ergodicity for classical kinetic SDEs
Condition (B) is equivalent to [43, Assumption  1], which can be seen by noting that $r_{0}\in(-1,1)$ and considering the matrix
@@
As detailed in [43, Example 1], for any $\gamma>0$, there exists $\kappa>0$ such that (B) is satisfied for $b(x,y)=-x-\gamma y+F(x,y)$ provided $|\nabla F|\leqslant\kappa$ outside a compact set.
@@
In the uniform dissipative case, [59] uses hypercontractivity to derive the exponential ergodicity in relative entropy. Let us remark that in the uniform dissipative case, the hypercontractivity can be derived by Wang’s Haranck inequality with power while in the partially dissipative case, Wang’s Harnack inequality can only derive hyperboundedness, see [28]. Recently, [43] derives the $L^{2}$-exponential ergodicity under the partially dissipative condition. Note that exponential ergodicity in relative entropy is strictly stronger than $L^{2}$-exponential ergodicity, see for instance [59, Proposition 2.3]. Together with hyperboundedness, $L^{2}$-exponential ergodicity implies exponential ergodicity in relative entropy, as we shall recall in Lemma 3.1.
@@ 3 Proof of the Main Theorems
Before we go on and prove our main results, we shall mention the following key result for our analysis. In order to do so, consider a conservative diffusion process on $\mathbb{R}^{m}$ with generator $L=\frac{1}{2}\mathrm{tr}(a\nabla^{2})+\langle b,\nabla\rangle,$ and associated semigroup $P_{t}$. Let $\mu$ be an invariant probability measure of $(P_{t})_{t\geq 0}$. By [59, Proposition 2.3], the exponential convergence for $\mu(P_{t}f\log P_{t}f)$ can be derived from hypercontractivity of $P_{t}$, i.e. $\|P_{t_{0}}\|_{L^{2}(\mu)\rightarrow L^{4}(\mu)}\leq 1$ for some $t_{0}>0$. Moreover, in that case, the entropy convergence rate depends only on $t_{0}$. On the other hand, [4, Step (d) in the proof of Theorem 1.1] states that hypercontractivity of $P_{t}$ can be deduced by the exponential convergence of $\mu((P_{t}f)^{2})$ together with hyperboundedness $\|P_{\tilde{t}_{0}}\|_{L^{2}(\mu)\rightarrow L^{4}(\mu)}\leq C$ for some constants $\tilde{t}_{0}>0,C>0$. Building on [59, Proposition 2.3] and [4, Step (d) in Proof of Theorem 1.1], we summarize the following key result on exponential ergodicity in relative entropy.
@@
We first present a lemma on the $L^{2}$-exponential ergodicity for (2.5). The proof largely follows the approach in [43, Theorem 2], where the technique of hypocoercivity plays a central role.
@@
(Step 1) The existence and uniqueness of an invariant probability measure is known, see for instance [28, Proof of Theorem 3.3]. By [28, Theorem 3.4], there exist constants $t_{0}>0$, $C_{0}>0$ such that
@@
Moreover, it follows from [28, Theorem 3.5] that the log-Sobolev inequality holds:
@@
for some $c_{1}>0$. By [61, Proposition 1.4.4(3)] this implies
@@
Under (C), (3.11) is well-posed and has a unique invariant probability measure $\Phi(\mu)\in\mathscr{P}_{2}(\mathbb{R}^{2d})$, see for instance [28, Theorem 3.3]. Let $\tilde{P}_{t}^{\mu}$ be the associated semigroup to (3.11). For any $M>0$, let
@@
We first give a log-Sobolev inequality for $\Phi(\mu)$ uniform over $\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d})$, which is a corollary of [28, Theorem 3.5(2)].
@@
By [28, Theorem 3.5(2)], it is sufficient to prove that for any $R>0$ and $M>0$, there exist positive constants $C_{1}$ and $C_{2}$ depending on $R$, $M$ such that
@@
Moreover, it follows from [28, (3.12)] that $(\Phi(\mu))_{\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d})}$ is tight so that there exists a constant $R_{0}>0$ such that
@@
(i) Thanks to [28, Theorem 3.3], there exists a constant $\eta_{0}>0$ such that when $K_{I}<\eta_{0}$, we can find a constant $M>0$ such that $\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d})$ is fixed by the map $\Phi$. By [28, Theorem 3.4], there exist constants $t_{0}>0$ and $C_{0}>0$ such that
@@
It now only remains to repeat the proof of [28, Theorem 3.9]. To be self-contained, we list the main procedure here.
@@
So, when $K_{I}$ is small enough (independently from $N$), it follows from [28, Proof of Theorem 3.3] that (2.19) has a unique invariant probability measure $\bar{\mu}^{N}\in\mathscr{P}_{2}((\mathbb{R}^{2d})^{N})$.
@@
From this, [61, Proposition 1.4.4(3)] gives
@@
In view of [28, (3.12)], there exists a constant $\varepsilon>0$ such that
@@
This together with [20, Theorem 1] implies

### v2

@@ 1 Introduction
The first exponential convergence result stated in terms of relative entropy has been obtained in the seminal work [50] of Villani. It introduced the notion of hypocoercivity which has been developed since then in various directions, see [17, 38, 1, 10, 6, 8] and references therein on this general topic. More specifically, [50, Theorem 39] gives an exponential decay in relative entropy along (1.3) provided $V\in C^{2}(\mathbb{R}^{d})$ and has a bounded Hessian, and $\mu$ satisfies a so-called log-Sobolev inequality (see (3.3) below) with respect to the Dirichlet form $\mathscr{E}(f,g)=\mu(\nabla f\cdot\nabla g)$, where $\nabla$ is the gradient on $\mathbb{R}^{2d}$.
@@
However, the result in [50, Theorem 39], along with nearly all subsequent works, relies on having an explicit representation of the invariant probability measure $\mu$. An exception arises in the so-called uniformly dissipative case, which is a very restrictive situation where the drift of the SDE induces a deterministic contraction along the stochastic flow [5, 35, 36]. In particular, there are many tools to establish log-Sobolev inequalities [3], but most of them require an explicit expression, or are restricted to the uniformly-dissipative case as the Bakry-Émery criterion (which in its classical form is written for explicit invariant measures but in fact apply more generally [36]).
@@ 2.1 Exponential Ergodicity for classical kinetic SDEs
Condition (B) is equivalent to [37, Assumption  1], which can be seen by noting that $r_{0}\in(-1,1)$ and considering the matrix
@@
As detailed in [37, Example 1], for any $\gamma>0$, there exists $\kappa>0$ such that (B) is satisfied for $b(x,y)=-x-\gamma y+F(x,y)$ provided $|\nabla F|\leqslant\kappa$ outside a compact set.
@@
In the uniform dissipative case, [51] uses hypercontractivity to derive the exponential ergodicity in relative entropy. Let us remark that in the uniform dissipative case, the hypercontractivity can be derived by Wang’s Haranck inequality with power while in the partially dissipative case, Wang’s Harnack inequality can only derive hyperboundedness, see [25]. Recently, [37] derives the $L^{2}$-exponential ergodicity under the partially dissipative condition. Note that exponential ergodicity in relative entropy is strictly stronger than $L^{2}$-exponential ergodicity, see for instance [51, Proposition 2.3]. Together with hyperboundedness, $L^{2}$-exponential ergodicity implies exponential ergodicity in relative entropy, as we shall recall in Lemma 3.1.
@@ 3 Proof of the Main Theorems
Before we go on and prove our main results, we shall mention the following key result for our analysis. In order to do so, consider a conservative diffusion process on $\mathbb{R}^{m}$ with generator $L=\frac{1}{2}\mathrm{tr}(a\nabla^{2})+\langle b,\nabla\rangle,$ and associated semigroup $P_{t}$. Let $\mu$ be an invariant probability measure of $(P_{t})_{t\geq 0}$. By [51, Proposition 2.3], the exponential convergence for $\mu(P_{t}f\log P_{t}f)$ can be derived from hypercontractivity of $P_{t}$, i.e. $\|P_{t_{0}}\|_{L^{2}(\mu)\rightarrow L^{4}(\mu)}\leq 1$ for some $t_{0}>0$. Moreover, in that case, the entropy convergence rate depends only on $t_{0}$. On the other hand, [4, Step (d) in the proof of Theorem 1.1] states that hypercontractivity of $P_{t}$ can be deduced by the exponential convergence of $\mu((P_{t}f)^{2})$ together with hyperboundedness $\|P_{\tilde{t}_{0}}\|_{L^{2}(\mu)\rightarrow L^{4}(\mu)}\leq C$ for some constants $\tilde{t}_{0}>0,C>0$. Building on [51, Proposition 2.3] and [4, Step (d) in Proof of Theorem 1.1], we summarize the following key result on exponential ergodicity in relative entropy.
@@
We first present a lemma on the $L^{2}$-exponential ergodicity for (2.5). The proof largely follows the approach in [37, Theorem 2], where the technique of hypocoercivity plays a central role.
@@
(Step 1) The existence and uniqueness of an invariant probability measure is known, see for instance [25, Proof of Theorem 3.3]. By [25, Theorem 3.4], there exist constants $t_{0}>0$, $C_{0}>0$ such that
@@
Moreover, it follows from [25, Theorem 3.5] that the log-Sobolev inequality holds:
@@
for some $c_{1}>0$. By [52, Proposition 1.4.4(3)] this implies
@@
Under (C), (3.11) is well-posed and has a unique invariant probability measure $\Phi(\mu)\in\mathscr{P}_{2}(\mathbb{R}^{2d})$, see for instance [25, Theorem 3.3]. Let $\tilde{P}_{t}^{\mu}$ be the associated semigroup to (3.11). For any $M>0$, let
@@
We first give a log-Sobolev inequality for $\Phi(\mu)$ uniform over $\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d})$, which is a corollary of [25, Theorem 3.5(2)].
@@
By [25, Theorem 3.5(2)], it is sufficient to prove that for any $R>0$ and $M>0$, there exist positive constants $C_{1}$ and $C_{2}$ depending on $R$, $M$ such that
@@
Moreover, it follows from [25, (3.12)] that $(\Phi(\mu))_{\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d})}$ is tight so that there exists a constant $R_{0}>0$ such that
@@
(i) Thanks to [25, Theorem 3.3], there exists a constant $\eta_{0}>0$ such that when $K_{I}<\eta_{0}$, we can find a constant $M>0$ such that $\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d})$ is fixed by the map $\Phi$. By [25, Theorem 3.4], there exist constants $t_{0}>0$ and $C_{0}>0$ such that
@@
It now only remains to repeat the proof of [25, Theorem 3.9]. To be self-contained, we list the main procedure here.
@@
So, when $K_{I}$ is small enough (independently from $N$), it follows from [25, Proof of Theorem 3.3] that (2.19) has a unique invariant probability measure $\bar{\mu}^{N}\in\mathscr{P}_{2}((\mathbb{R}^{2d})^{N})$.
@@
From this, [52, Proposition 1.4.4(3)] gives
@@
In view of [25, (3.12)], there exists a constant $\varepsilon>0$ such that
@@
This together with [19, Theorem 1] implies